### v2

Abstract
We show that there exists an infinite family of cubic 222-connected non-hamiltonian graphs with girth 555 containing a unique longest cycle.
MSC 2020: 05C45, 05C38
Keywords:
Longest cycle Girth Connectivity
1 Introduction
The result proven in this note is fundamentally motivated by the problem of counting hamiltonian cycles in regular graphs, with a focus on hamiltonian regular graphs with few hamiltonian cycles. Perhaps the most important problem in this area is Sheehan’s Conjecture [10], stating that every hamiltonian 4-regular graph has at least two distinct hamiltonian cycles. For recent articles treating the problem theoretically, see [1, 5, 6, 7, 14], and for a recent algorithmic treatment, see [3]. In this context we also want to highlight an interesting recent conjecture by Girão, Kittipassorn and Narayanan [5] stating that there exists an absolute constant $K>0$ such that if an $n$-vertex graph $G$ with minimum degree at least 333 contains a hamiltonian cycle, then $G$ contains another cycle of length at least $n-K$.
It has been known since the forties that if a 3-regular graph contains one hamiltonian cycle, then it must contain at least three [12]. Chia and Thomassen explored in [2] a natural variation of this problem: instead of requiring the graph to contain a unique hamiltonian cycle, they investigated graphs with a unique longest cycle. Although 3-regular graphs with a unique hamiltonian cycle do not exist, it is not difficult to describe 3-regular graphs containing exactly one longest cycle. For examples of such graphs satisfying various additional properties—e.g. regarding connectivity, girth, and length of the longest cycle—see for instance [2] itself and [14].
Inspired by an old conjecture of Cantoni [13] positing that every planar 3-regular graph with exactly three hamiltonian cycles contains a triangle, a question raised in [14] by the second author asks whether 3-regular 2-connected triangle-free graphs containing a unique longest cycle exist. Motivated, in turn, by this question, the first author posed a relaxation of this problem [7] asking for the minimum number of vertices with a degree different from 3 amongst all 2-connected triangle-free non-hamiltonian graphs with minimum degree 3 containing a unique longest cycle. Moreover, he presented an example of a 3-regular 2-connected triangle-free non-hamiltonian graph containing precisely four longest cycles as well as a 2-connected triangle-free non-hamiltonian 60-vertex graph having a unique longest cycle and 56 vertices of degree 3 and 4 vertices of degree 4. In this note we answer in the affirmative the initial question, and thus also solve the proposed relaxation.
2 The result
Theorem. If there exists a cubic $\kappa_{1}$-connected graph $G_{1}$ with girth $g_{1}$ having a unique longest cycle, a cubic graph $G_{2}$ with girth $g_{2}$ having precisely three hamiltonian cycles, and a vertex $u\in V(G_{2})$ such that $G_{2}-u$ is $\kappa_{2}$-connected, then there exists an infinite family of cubic $\min\{\kappa_{1},\kappa_{2}\}$-connected graphs with girth at least $\min\{2g_{1},g_{2}\}$ having a unique longest cycle.
Proof
Let $v,w$ and $x$ be the three neighbors of $u$ in $G_{2}$. For a graph $G$ and a vertex $a\in V(G)$ with precisely three neighbors $b$, $c$ and $d$, we define the marriage of vertex $a\in V(G)$ with $(G_{2},u)$ as one of the graphs obtained by taking the disjoint union of $G-a$ and $G_{2}-u$ and adding three edges such that the resulting graph is cubic and every added edge has one endpoint in $\{b,c,d\}$ and the other endpoint in $\{v,w,x\}$ (for the purpose of this proof, it does not matter which of the six possible graphs we choose). Kotzig and Labelle [8] call this operation marriage, whereas for example Eiben, Jajcay and Šparl [4] refer to a very similar operation as generalized truncation. For each vertex $t\in V(G_{1})$, let $G_{2,t}$ be a copy of $G_{2}$ and let $u_{t}\in V(G_{2,t})$ be the vertex corresponding to $u\in V(G_{2})$. Let $H$ be a graph obtained by marrying every vertex $t\in V(G_{1}$) with $(G_{2,t},u_{t})$. We label the vertices of $H$ as $s_{i,j}$ in the natural way ($i\in V(G_{1}),j\in V(G_{2}-u)$).
$H$ is clearly cubic by construction. It is easy to check that for arbitrary vertices $p\in V(G_{1})$ and $q\in V(G_{2})$ if $G_{1}$ is $\kappa_{1}$-connected and $G_{2}-q$ is $\kappa_{2}$-connected, then the marriage of $p$ and $(G_{2},q)$ is at least $\min\{\kappa_{1},\kappa_{2}\}$-connected. Therefore $H$ is also $\min\{\kappa_{1},\kappa_{2}\}$-connected. Let $\mathcal{C}=s_{i_{0},j_{0}}s_{i_{1},j_{1}}\ldots s_{i_{m-1},j_{m-1}}s_{i_{0},j_{0}}$ be any cycle of length $m$ in $H$. If $i_{0}=i_{1}=\ldots=i_{m-1}$, then $j_{0}j_{1}\ldots j_{m-1}j_{0}$ is a cycle of length $m$ in $G_{2}-u$. If $i_{k}\neq i_{k+1}$ for some $0\leq k\leq m-1$, then $i_{k-1}=i_{k}$ and $i_{k+1}=i_{k+2}$ (where indices are taken modulo $m$), because $H$ is obtained by marrying the vertices of $G_{1}$ with $(G_{2},u)$. Starting from the index $i_{k+1}$, generate a new sequence of indices $\mathcal{C}_{2}$ by moving sequentially through the indices, wrapping around from $i_{m-1}$ to $i_{0}$, until $i_{k+1}$ is reached again. At each step, include an index in $\mathcal{C}_{2}$ only if it is different from the immediately preceding index. Now $\mathcal{C}_{2}$ corresponds to a cycle in $G_{1}$. Hence, every cycle in $H$ corresponds to either a cycle in $G_{2}-u$ or a cycle in $G_{1}$ in which every vertex on that cycle is replaced by a path of length at least one. Clearly, the girth of $G_{2}-u$ is at least $g_{2}$. Therefore, the girth of $H$ is at least $\min\{2g_{1},g_{2}\}$.
Thomason [11] showed that in any graph where all vertices have an odd degree every edge is contained in an even number of hamiltonian cycles. Because of Thomason’s result, $G_{2}$ contains precisely one hamiltonian cycle containing simultaneously the edges $uv$ and $uw$, one hamiltonian cycle containing simultaneously the edges $uv$ and $ux$ and one hamiltonian cycle containing simultaneously the edges $uw$ and $ux$. As a result, the graph $G_{2}-u$ has precisely one hamiltonian $vw$-path, one hamiltonian $vx$-path and one hamiltonian $wx$-path. Hence, if $\mathcal{C}_{1}$ is the unique longest cycle in $G_{1}$, then the longest cycle in $H$ can have length at most $|\mathcal{C}_{1}|\cdot|V(G_{2}-u)|$ and there is indeed such a cycle (obtained by replacing each vertex in $\mathcal{C}_{1}$ by an appropriate hamiltonian path in $G_{2}-u$). Conversely, by contracting in $H$ every copy of $G_{2}-u$ to a single vertex, every longest cycle in $H$ corresponds to a longest cycle in $G_{1}$. Since $G_{2}-u$ has a unique hamiltonian $yz$-path for $y,z\in\{v,w,x\},y\neq z$, we conclude that $H$ has a unique longest cycle.
By iteratively applying this argument, where $G_{1}$ is replaced by $H$, we obtain the desired infinite family.
We remark that every cubic graph containing a unique longest cycle is necessarily non-hamiltonian due to Smith’s observation in [12]. In fact more strongly, one can infer that the length of the unique longest cycle in such a graph is at most $n-2$, where $n$ is the order of the graph, because Thomason [11] proved that for every graph $G$ in which all vertices have an odd degree, the number of hamiltonian cycles in $G$ has the same parity as the number of hamiltonian cycles in $G-v$ for all $v\in V(G)$.
Chia and Thomassen [2] constructed a cubic 2-connected graph $G_{1}$ with girth 3 having a unique longest cycle (shown in Fig. 1). Schwenk [9] constructed a cubic 3-connected graph $G_{2}$ with girth 5 having precisely three hamiltonian cycles (shown in Fig. 2). Therefore, for every vertex $u\in V(G_{2})$, the graph $G_{2}-u$ is 2-connected. Note that if $G$ is a graph with girth at least 555, then for any vertex $u\in V(G)$ the graph $G-u$ has girth at least 555, but for our specific graph $G_{2}$, the graph $G_{2}-u$ has girth exactly 555. Based on this observation and by applying the previous theorem, we obtain the following corollary.
Figure 1: A cubic 2-connected graph with girth 3 having a unique longest cycle.
Figure 2: A cubic 3-connected graph with girth 5 having precisely three hamiltonian cycles (this graph is known as the generalised Petersen graph GP(9,2)).
Corollary. There exists an infinite family of cubic 222-connected non-hamiltonian graphs with girth 555 containing a unique longest cycle.
3 Two open problems
We conclude this note with two questions that spark our interest.
Question 1
Does there exist a cubic 222-connected non-hamiltonian planar triangle-free graph containing a unique longest cycle?
If Cantoni’s conjecture [13] holds, it seems that a technique different from the one presented in the theorem will be necessary to construct such graphs if they exist.
The second question that interests us is as follows:
Question 2
Does there exist a cubic 333-connected graph with girth at least $g$ having precisely three hamiltonian cycles for each integer $g$?
If the answer is positive, one can infer the existence of an infinite family of cubic 2-connected non-hamiltonian graphs with girth at least $g$ containing a unique longest cycle for each integer $g$ (by using the corollary and then iteratively applying the theorem).
Acknowledgements
Jorik Jooken is supported by a Postdoctoral Fellowship of the Research Foundation Flanders (FWO) with grant number 1222524N.